OpenStudy (anonymous):

In a basketball league, the Falcons won 15 of their first 20 games. In order to win 60% of the total number of games, how many more games will they have to play, assuming they win only half of the remaining games?

OpenStudy (anonymous):

here is what I have but it don't look to be good help me

OpenStudy (anonymous):

15 + 0.5x = 0.6(20 + x) 15 + 0.5x = 12 + 0.6x 3 = 0.1x x = 3/0.1 = 30. The Falcons would have to play 30 additional games.

OpenStudy (anonymous):

and she ask me this How many games is 0.5x?